OpenStudy (anonymous):

antiderivative. Given f'''(x)=e^x with f''(0)=6, f'(0)=5, then f(x)=______+C

OpenStudy (anonymous):

Integrate, sub for the constant, repeat....

OpenStudy (anonymous):

\[\int e^xdx =e^x+C\] \[f''(x)=e^x+C\] and \[f''(0)=e^0+C=6\] \[1+C=6\] \[C=5\] so \[f''(x)=e^x+5\] \[f'(x)=\int e^x+5 dx = e^x+5x+C\]

OpenStudy (anonymous):

\[f'(0)=5=e^0+5\times 0+C=1+C\] \[C=4\] so \[f'(x)=e^x+5x+4\]

OpenStudy (anonymous):

one more time and you get your answer